OpenStudy (anonymous):

4x^2 + 9y^2 =72 x^2 - y^2 = 5 systems of equations

hero (hero):

One easy way to solve this is to perform the following steps: Multiply both sides of the second equation by 4. Doing that will yield the following system: 4x^2 + 9y^2 = 72 4x^2 - 4y^2 = 20 Subtracting the second equation from the first yields: 13y^2 = 52 Then divide both sides by 13: y^2 = 4 y = + or - 2 To find x, you can use substitution or you can do this: Multiply the second equation by 9 to get this system: 4x^2 + 9y^2 = 72 9x^2 - 9y^2 = 45 Combine both equations to get: 13x^2 = 117 Divide both sides by 13 to get x^2 = 9 x = + or - 3 Now, that yields 4 possible points: (-3, -2), (3, -2), (-3, 2) and (3, 2)

hero (hero):

@Piercethesyd

OpenStudy (anonymous):

Will these always have 4 answers unless either x or y is 0?

hero (hero):

Not always. Take the following system for example: http://www.wolframalpha.com/input/?i=x%5E2+%2B+y%5E2+%3D+5+and+x%5E2+%2B+y%5E2+%3D+4

hero (hero):

There are no intersections for that particular system.

OpenStudy (anonymous):

Cool, like the equivalent of parallel lines in linear systems of equations.

hero (hero):

lol, yes

hero (hero):

It's actually called "concentric" rather than "parallel" in this case.

OpenStudy (anonymous):

Okay, I understand. I was just drawing an equivalent, but the term is helpful.

hero (hero):

As long as you're learning, that's all that matters.

hero (hero):

Here's a system that yields only two solutions: http://www.wolframalpha.com/input/?i=x%5E2+-+y%5E2+%3D+5+and+5x%5E2+%2B+6y%5E2+%3D+25